A glider is traveling 240 kilometers in 3 hours with the wind. It travels 210 kilometers in 3 hours against the wind. What is the speed of the wind and the glider?
Answer by ankor@dixie-net.com(12701)

A glider is traveling 240 kilometers in 3 hours with the wind. It travels 210 kilometers in 3 hours against the wind. What is the speed of the wind and the glider?
:
How does a glider go against the wind?
:
But assuming it is powered somehow, here is the solution:
:
Let x = speed of the plane
Let y = speed of the wind:
Then
(x+y) = ground speed with the wind
and
(x-y) = ground speed against the wind
:
Write two distance equations; Distance = time * speed
:
3(x+y) = 240; with the wind
3(x-y) = 210: against the wind
:
We can simplify both equations by dividing them by 3, leaving us with:
x + y = 80
x - y = 70
------------adding eliminates y; find x
2x + 0 = 150
x = 150/2
x = 75 km/hr speed of the "glider"
:
Find the wind speed
75 + y = 80
y = 80 - 75
y = 5 km/hr is the wind
:
:
Check solutions in one of the original equations
3(75 - 5) =
3(70) = 210 km
:
Did this make sense? Any questions?